OpenStudy (anonymous):

how much energy is needed to completely remove an electron from n=4?

OpenStudy (praxer):

Energy required is \[E_\infty -E_4\] \[E_ \infty=0\] \[E_4=13.6/4^2\] therefore E= Do the calculation.

OpenStudy (praxer):

\[E_n=13.6/n^2\]

OpenStudy (praxer):

@talita did you understand this.

OpenStudy (anonymous):

@praxer That only works for hydrogen :P

OpenStudy (ipwnbunnies):

^^ This is much important. It would depend on the electron config of the atom.